OpenStudy (anonymous):

What is the missing number? (–5.7 • 2) • 8 = –5.7 • (2 • ? )

OpenStudy (anonymous):

\[(-5.7\times2)\times8=-5.7\times(2\times?)\]

OpenStudy (ddcamp):

\[Associative\ Property:\\ (a * b) * c = a * (b * c)\]

OpenStudy (anonymous):

i know that

OpenStudy (anonymous):

what is the missing number to my question

OpenStudy (ddcamp):

For your problem, what are a, b, and c?

OpenStudy (anonymous):

omg thank you so much

OpenStudy (anonymous):

\[(−5.7×2)×8=−5.7×(2×8)\]

OpenStudy (ddcamp):

Yup!